A garment factory distributes two brands of jeans. If it is found that 75 out of 250 customers prefer brand A and that 30 out of 150 prefer brand B, can we conclude at 0.05 level of significance that brand A outsells brand B?

1
Expert's answer
2022-05-31T06:54:31-0400

a.

"H_0: p_1\le p_2"

"H_a:p_1>p_2"

b.The significance level is "\alpha = 0.05."

c. This corresponds to a right-tailed test, and a z-test for two population proportions will be used.

d. The value of the pooled proportion is computed as



"\bar{p}=\dfrac{X_1+X_2}{n_1+n_2}=\dfrac{75+30}{250+150}=0.2625"

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}_1-\hat{p}_2}{\sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}""=\dfrac{75/250+30/105}{\sqrt{0.2625(1-0.2625)(1/250+1/150)}}=2.2006"

e. Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

f. The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

g. Since it is observed that "z = 2.2006 >1.6449= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.2006)=0.013882," and since "p=0.013882<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p_1" is greater than "p_2," at the "\alpha = 0.05" significance level.